OpenStudy (anonymous):

Simplify (x^2 + 16)(x^2 – 16) = ? Btw, i don't want a direct answer. I want an explanation as to how to achieve a correct answer. (P.s.) Thanks a ton! ^_^

OpenStudy (anonymous):

@AravindG @pooja195

OpenStudy (anonymous):

@dan815 @acxbox22

jhonyy9 (jhonyy9):

do yu know fomula a^2 -b^2 = (a-b)(a+b) yes ? try us this here --- ok ?

OpenStudy (anonymous):

k, thanks. I'll try it. ^_^

jhonyy9 (jhonyy9):

i like see it now - ok ?

OpenStudy (anonymous):

so, basically, i do: x^2^2 - 16^2 ?

jhonyy9 (jhonyy9):

yes but you need simplify it not make more complex

OpenStudy (anonymous):

so, i do: x^3 - 16^2 ?

jhonyy9 (jhonyy9):

do you understand my words above ?

OpenStudy (anonymous):

i think so?

jhonyy9 (jhonyy9):

ok than try it to simplify - ok ?

OpenStudy (anonymous):

so... x^4 + 32?

jhonyy9 (jhonyy9):

how you have got 32 ? 16*16 not equal 32 - yes ? so but i have said to you above that you need simplify it not make more complex - yes ?

OpenStudy (anonymous):

oh, sorry i mathed wrongly. i meant: x^4 + 256

OpenStudy (anonymous):

that one is it right?

jhonyy9 (jhonyy9):

do you have same choices ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

here are my choices A. x^4 – 256 B. x^4 + 32 C. x^4 + 256 D. x^4 – 32

jhonyy9 (jhonyy9):

but this not is right because formula say that (a-b)(a+b)=a^2 -b^2

jhonyy9 (jhonyy9):

do you see where is wrong your above answer ?

OpenStudy (anonymous):

so ... i do A. than?

OpenStudy (anonymous):

x^4 - 256?

jhonyy9 (jhonyy9):

yes sure

OpenStudy (anonymous):

Thanks! ^_^

jhonyy9 (jhonyy9):

was my pleasure - so do you understand it why ?

OpenStudy (anonymous):

yes, i understand now, i was not supposed to use +, because it was -

jhonyy9 (jhonyy9):

yes but you need learn formulas beaucose you need us them always

OpenStudy (anonymous):

Thanks again! ^_^